# Enabling data-driven and bidirectional model development in Verilog-A for photonic devices

### v2

@@
journal: opticajournalarticletype: Research Article{abstract*}
@@ 2.2 Implementation
where $b_{i}$ is the outgoing wave at the port $i$ and $a_{j}$ is the incoming wave at the port $j$. The introduction of s-parameters is also advantageous for data-driven model development since all the photonic blocks can be characterized by the frequency dependent s-parameter data. We note that the impedance at which these devices are measured is not necessarily related to the reference impedance $R_{R}$. The latter is used as a knob to control the reflectivity of a component from each port in the absence of measurement data. In Verilog-A, one can import s-parameters as a Lookup Table (LUT) and write the equations for the outgoing waves using the s-parameters. For a generic n-port network this can mathematically be described as:
@@ 3 Simulation
To demonstrate the efficacy of our modeling approach, we built a Michelson Interferometer (MI) using a Verilog-A library of silicon photonic devices [5, 7]. As per the typical MI design, it consists of 3-dB directional coupler, waveguides, and loop mirrors built from Y-branch and waveguide bends. Note that the laser is internally set to have an isolator in order to keep feeding constant power, though this can be modified in the code. Passive photonic device models in the library are built using the data from [23]. Active components are developed through a combination of analytical description and simulation/measurement data [24, 25]. Laser is the only device in our library that has purely behavioural description because we use it as a continuous wave (CW) supply to the PICs. Fig. 2a depicts the schematic view of the MI assembled in Cadence Virtuoso [3]. The corresponding layout of the fabricated device is shown in Fig. 2b. MI is an ideal testbench for demonstration of bidirectional signaling since the forward and backward propagating signals are simultaneously present in each arm. Furthermore, signals in both arm should constructively interfere to be able to see the signal at the photodetector side. In case of destructive interference, all of the returning signals should appear at the laser side. In contrast to the example from [11], our testbench can be viewed to have three optical sources (one forward propagating signal from the laser and two reflected signals from the mirror loops) all meeting at the directional coupler. Thus, it does not have the limitation of the prior work. Fig. 3 compares the transmission response of the MI that has been measured against the Verilog-A based simulation. For each case the photonic device models are fitted to the measurements. One simulation was run using the bidirectional signaling methodology introduced above whereas the other was executed without this feature. It is clear that neglecting the bidirectional signal propagation leads to erroneous results.
@@
For transient analysis, we replace the waveguides in each arm of the previous example with the thermal and carrier-depletion phase shifters. We also add a transistor-level electronic driving circuit from the 65 nm CMOS technology library to modulate the carrier-depletion phase shifter. Fig. 4 shows the schematic view of the MIM simulated in the same environment as the CMOS driver. The detailed schematic of the CMOS driver circuit is displayed in Fig. 5. It is based on pseudo-differential inverters topology [26], which can provide 4.8V differential output voltage swing. The voltage swing requirements are dictated by the pn-junction phase shifter length, which is 800 $\mu m$ in this design. M1, M3, M5, and M7 are n-type MOSFETs, whereas M2, M4, M6, and M8 are p-type MOSFETs. The input signal arrangement is also shown in the same figure. The basic idea here is to alternately switch on a set of transistors (either M1, M3, M6, M8 or M2, M4, M5, M7) to create 0V-4.8V modulation voltage across the pn-junction diode, which, in its turn, changes the effective index of the doped waveguide, thereby modulating the amount of phase shift applied to the optical signal. The transient behavior of the circuit is provided in Fig. 6. The reflected signals arrive at the laser and photodetector interfaces in accordance with the modulation signal. Furthermore, as expected, the laser’s output signal (i.e. forward propagating signal) remains the same regardless of the incoming signals. Similarly, the photodetector’s forward propagating signal is zero at all times.
@@
Next, we consider a testbench for the Bragg gratings based Fabry-Perot Cavity (FPC) simulation, similar to what is presented in [18], wherein the Verilog-A photonic models were deemed incapable of simulating bidirectional signal propagation. Fig. 7 illustrates the developed testbench (Fig. 7a) and the layout of the corresponding fabricated device (Fig. 7b), where we designed the FPC to have a resonant peak at the wavelength close to 1541 nm. Since FPC is a circuit consisting of passive elements only, its baseband time domain response is static. Therefore the main focus is on its frequency spectrum. We run wavelength sweep simulation on the presented circuit and compare it against the measured FPC response to arrive at Fig. 8. The difference between the bidirectionality enabled simulation and conventional simulation is apparent. The latter fails to capture the bandpass behaviour at around 1541 nm, which may be critical for some applications. Similar to the previous testbench, the Verilog-A models here are also fitted to the measurement results. In this way, the proposed Verilog-A modeling fully captures the true FPC response, in contrast to the limitation described in [18].
@@
The final simulation scenario involves the creation of the testbench for photonic Reflection Cancellation Circuit (RCC), illustrated in Fig. 9a. If a fraction of the optical power emitted from the laser reflects back into the laser, it can lead to the broadening of its linewidth, increased Relative Intensity Noise (RIN), and eventually render the laser unstable. An RCC is a compact on-chip solution that ensures the stability of the laser by actively canceling out any reflections to the laser through its phase tunable reflector arm and the interference effect [13]. More details of this circuit can be found in [13]. It consists of: 1) 90/10 power tap connecting to the photodetector to track the reflection cancellation amount; 2) intermediate adjustable splitter ($V_{ps1}$) along with phase tunable reflector ($V_{ps2}$) to control the magnitude and phase of the reflection canceling signal; and 3) output stage adjustable splitter ($V_{att}$) used to control the amount of reflection coming back from the subsequent device. Fig. 9b demonstrates the layout, which, besides the RCC itself, contains routing and reflection generator stages (highlighted as 1 and 5 in the figure). While the output portion of the RCC is only used for the initial setup, the former two are critical for an effective cancellation of any reflection coming from the device. We use the $V_{ps1}$ and $V_{ps2}$ knobs to run 2D sweep of the reflection landscape and observe the optical signal strength at the input of the photodetector. Fig. 10 shows the resulting measurement and bidirectional signaling-enabled simulation response. The simulation results closely follow the real RCC behaviour, including the skewed response which comes from the imbalanced 3-dB directional couplers, and the asymmetric power requirements to reach the global minima which is the evidence of thermal crosstalk between the phase shifters [27]. We note that without the bidirectional signaling it would not be possible to capture local or global minima, which is essential for RCC’s functionality.

### v3

@@
journal: opticajournalarticletype: Research Article
{abstract*}
@@ 2.2 Implementation
where $b_{i}$ is the outgoing wave at the port $i$ and $a_{j}$ is the incoming wave at the port $j$. The introduction of s-parameters is also advantageous for data-driven model development since all the photonic blocks can be characterized by the frequency dependent s-parameter data. We note that the impedance at which these devices are measured is not necessarily related to the reference impedance $R_{R}$. The latter is used as a knob to control the reflectivity of a component from each port in the absence of measurement data. In Verilog-A, one can import s-parameters as a Lookup Table (LUT) and write the equations for the outgoing waves using the s-parameters [23]. For a generic n-port network this can mathematically be described as:
@@ 3 Simulation
3.1 Michelson Interferometer testbench
To demonstrate the efficacy of our modeling approach, we built a Michelson Interferometer (MI) using a Verilog-A library of silicon photonic devices [5, 7]. As per the typical MI design, it consists of 3-dB directional coupler, waveguides, and loop mirrors built from Y-branch and waveguide bends. Note that the laser is internally set to have an isolator in order to keep feeding constant power, though this can be modified in the code. In our library, all the passive photonic device models except the waveguide employ the data-driven approach and are built using the data from [24]. The waveguide is described using analytical equations. Active components such as photodetectors and modulators are developed through a combination of analytical description, curve-fits, and simulation/measurement data [5, 25, 26]. Thermal phase shifters are based on the analytical equations presented in [7]. Laser is the only device in our library that has purely behavioural description because we use it as a continuous wave (CW) supply to the PICs [5]. Fig. 2a depicts the schematic view of the MI assembled in Cadence Virtuoso [3]. The corresponding layout of the fabricated device is shown in Fig. 2b. MI is an ideal testbench for demonstration of bidirectional signaling since the forward and backward propagating signals are simultaneously present in each arm. Furthermore, signals in both arm should constructively interfere to be able to see the signal at the photodetector side. In case of destructive interference, all of the returning signals should appear at the laser side. In contrast to the example from [11], our testbench can be viewed to have three optical sources (one forward propagating signal from the laser and two reflected signals from the mirror loops) all meeting at the directional coupler. Thus, it does not have the limitation of the prior work. Fig. 3 compares the transmission response of the MI that has been measured against the Verilog-A based simulation. The simulated frequency spectrum is obtained by passing the wavelength as a parameter to each component and running wavelength sweep simulation. Also, in order to focus on the actual device response, we remove the grating coupler effects with the help of de-embedding structures. Following this, we adjust the waveguide effective index and loss to match the measurements. One simulation was run using the bidirectional signaling methodology introduced above whereas the other was executed without this feature. It is clear that neglecting the bidirectional signal propagation leads to erroneous results.
@@ 3.1 Michelson Interferometer testbench
3.2 Michelson Interferometer Modulator testbench
For transient analysis, we replace the waveguides in each arm of the previous example with the thermal and carrier-depletion phase shifters. We also add a transistor-level electronic driving circuit from the 65 nm CMOS technology library to modulate the carrier-depletion phase shifter. Fig. 4 shows the schematic view of the Michelson Interferometer Modulator (MIM) simulated in the same environment as the CMOS driver. The detailed schematic of the CMOS driver circuit is displayed in Fig. 5. It is based on pseudo-differential inverters topology [27], which can provide 4.8V differential output voltage swing. The voltage swing requirements are dictated by the pn-junction phase shifter length, which is 800 $\mu m$ in this design. M1, M3, M5, and M7 are n-type MOSFETs, whereas M2, M4, M6, and M8 are p-type MOSFETs. The input signal arrangement is also shown in the same figure. The basic idea here is to alternately switch on a set of transistors (either M1, M3, M6, M8 or M2, M4, M5, M7) to create 0V-4.8V modulation voltage across the pn-junction diode, which, in its turn, changes the effective index of the doped waveguide, thereby modulating the amount of phase shift applied to the optical signal. The transient behavior of the circuit is provided in Fig. 6. The reflected signals arrive at the laser and photodetector interfaces in accordance with the modulation signal. Furthermore, as expected, the laser’s output signal (i.e. forward propagating signal) remains the same regardless of the incoming signals. Similarly, the photodetector’s forward propagating signal is zero at all times.
@@ 3.3 Fabry-Perot Cavity testbench
Next, we consider a testbench for the Bragg gratings based Fabry-Perot Cavity (FPC) simulation, similar to what is presented in [18], wherein the Verilog-A photonic models were deemed incapable of simulating bidirectional signal propagation. Fig. 7 illustrates the developed testbench (Fig. 7a) and the layout of the corresponding fabricated device (Fig. 7b), where we designed the FPC to have a resonant peak at the wavelength close to 1541 nm. The fabricated device has 120 corrugations per reflector, each with a width of 50 nm, whereas the cavity length is equal to one Bragg period (317 nm). Since FPC is a circuit consisting of passive elements only, its baseband time domain response is static. Therefore the main focus is on its frequency spectrum. We run wavelength sweep simulation on the presented circuit and compare it against the measured FPC response to arrive at Fig. 8. The difference between the bidirectionality enabled simulation and conventional simulation is apparent. The latter fails to capture the bandpass behaviour at around 1541 nm, which may be critical for some applications. Similar to the previous testbench, the grating coupler effect is de-embedded from the actual device response and the Verilog-A models here are also fitted to the measurement results. In this way, the proposed Verilog-A modeling fully captures the true FPC response, in contrast to the limitation described in [18].
3.4 Reflection Cancellation Circuit testbench
@@
The final simulation scenario involves the creation of the testbench for photonic Reflection Cancellation Circuit (RCC), illustrated in Fig. 9a. If a fraction of the optical power emitted from the laser reflects back into the laser, it can lead to the broadening of its linewidth, increased Relative Intensity Noise (RIN), and eventually render the laser unstable. An RCC is a compact on-chip solution that ensures the stability of the laser by actively canceling out any reflections to the laser through its phase tunable reflector arm and the interference effect [13]. More details of this circuit can be found in [13]. It consists of: 1) 90/10 power tap connecting to the photodetector to track the reflection cancellation amount; 2) intermediate adjustable splitter ($V_{ps1}$) along with phase tunable reflector ($V_{ps2}$) to control the magnitude and phase of the reflection canceling signal; and 3) output stage adjustable splitter ($V_{att}$) used to control the amount of reflection coming back from the subsequent device. Fig. 9b demonstrates the layout, which, besides the RCC itself, contains routing and reflection generator stages (highlighted as 1 and 5 in the figure). While the output portion of the RCC is only used for the initial setup, the former two are critical for an effective cancellation of any reflection coming from the device. We use the $V_{ps1}$ and $V_{ps2}$ knobs to run 2D sweep of the reflection landscape and observe the optical signal strength at the input of the photodetector. Fig. 10 shows the resulting measurement and bidirectional signaling-enabled simulation response. The simulation results closely follow the real RCC behaviour, including the skewed response which comes from the imbalanced 3-dB directional couplers, and the asymmetric power requirements to reach the global minima which is the evidence of thermal crosstalk between the phase shifters [28]. The thermal crosstalk is modeled by introducing a crosstalk parameter into the voltage source of each thermal phase shifter. Note that in the simulations we do not include grating couplers and therefore achieve higher optical power readings. Additionally, one can observe deviations in the electrical power requirements necessary to arrive at a global minima between the two plots. This could be attributed to the fact that our thermal phase shifter model is purely analytical and may have different thermal characteristics. We emphasize that without the bidirectional signaling it would not be possible to capture local or global minima, which is essential for RCC’s functionality.